A car is traveling at a speed of 30 m/s. What is its speed in kilometers per hour? CF: 1 km = 10^3 m, 1 h = 3600 s *
1
Expert's answer
2020-10-22T18:08:12-0400

Let's find the speed of the car in kilometers per hour:


"v=(30\ \dfrac{m}{s})\cdot (\dfrac{1\ km}{10^3 m})\cdot(\dfrac{3600\ s}{1\ h})=108\ \dfrac{km}{h}."

Answer:

"v=108\ \dfrac{km}{h}."